Circular No. 6325

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

NOVALIKE VARIABLE IN SAGITTARIUS S. Benetti, H. W. Duerbeck, and W. C. Seitter, European Southern

Observatory; and T. Harrison, New Mexico State University, write: “CCD spectrograms (439–702 nm, resolution 0.28 nm) taken of this object (IAUC 6322) by W. Hoff (Astrophysikalisches Institut Jena) on Feb. 24.4 and 25.4 UT with the ESO 1.52-m telescope (+ Boller & Chivens spectrograph) show a pure absorption-line spectrum with lines of He I, C I, C II, N I, O I, Si II, and faint lines of H. The lines are resolved (FWHM 90– 100 km/s), the average heliocentric radial velocity is +108 km/s, and the strong interstellar Na I D lines have a radial velocity of +3 km/s. The slow brightness evolution, and the C-rich and H-poor spectrum, make this object the first candidate for a star undergoing its final helium flash since the eruption of V605 Aql in 1919 (cf. Seitter 1989, IAU Symp. 131, p. 315).”

SUPERNOVA 1996C IN MCG +8-25-47 P. Sicoli reports the following precise position of SN 1996C, obtained

on Feb. 22.994 UT by A. Testa and P. Ghezzi at Sormano with the 0.5-m reflector (+ CCD): α = 13^{h}50^{m}48^{s}.60, δ = +49^{o}19′07′′.1 (equinox 2000.0).

D. Balam, University of Victoria, gives the following position end figures for SN 1996C: 48^{s}.63, 08′′.3. The supernova is located 13′′.4 north and 1′′.6 west of the galaxy’s nucleus; a nearby star (V = 14.6) has position end figures 47^{s}.80, 18′38′′.9. Images (Johnson V) obtained with the 1.82-m Plaskett reflector of the Dominion Astrophysical Observatory yielded the following preliminary magnitudes: Feb. 24.47 UT, 16.4; 25.40, 16.4.

GK PERSEI J. A. Mattei, AAVSO, reports that this old nova (Nova Per 1901) has

gone into a minor outburst for the first time since 1992 (when it reached mag 10.3 in late July; cf. IAUC 5558), as indicated by the following observations reported to the AAVSO: Feb. 19.03 UT, 12.8 (J. Bortle, Stormville, NY); 20.09, 12.9 (W. Dillon, Missouri City, TX); 23.13, 13.0 (R. Royer, Lakewood, CA); 23.83, 12.9 (P. Schmeer, Saarbrücken, Germany); 25.05, 12.7 (R. A. Komorous, London, ON); 25.87, 12.6 (J. R. Osorio, Madrid, Spain); 26.05, 12.4 (J. McKenna, Upper Montclair, NJ).

1996 February 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green